Time to grow: You are able to deposit $850 in a bank CD today, and you will withdraw the money only once the balance is $1,000. If the bank pays 5 percent interest, how long will it take you to attain your goal?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1000=850*(1.05)^n

(1000/850)=(1.05)^n

Taking log on both sides;

log (1000/850)=n*log 1.05

n=log (1000/850)/log 1.05

=3.33 years(Approx).